Suppose the demand curve for a product is given by Q = 300 - 2P + 4I, where I is average income measured in thousands of dollars. The supply curve is Q = 3P - 50. If I = 25, find the market-clearing price and quantity for the product.


1
Expert's answer
2021-10-18T18:09:51-0400



Given that I=25, the demand curve will be

Q=300-2p+(4*25)

Q=300-2P+100, Thus Q=400-2P

Supply curve is Q=3P-50

For market price and quantity, demand=supply

400-2P=3P-50

5P=450, Thus, P=90

400-(2*90) =220

Therefore, the market price is 90 and the quantity is 220